OpenStudy (anonymous):

evaluate exponential expression.. 5^3 x 5^5

jimthompson5910 (jim_thompson5910):

When multiplying exponential expressions, you add the exponents \[\Large 5^3 \times 5^5 = 5^{3+5}\]

OpenStudy (anonymous):

2^-3 x 2 ?

jimthompson5910 (jim_thompson5910):

2 = 2^1 Add the exponents and use the idea that \[\Large x^{-k} = \frac{1}{x^k}\] where k is some positive number

jimthompson5910 (jim_thompson5910):

Ex: \[\Large 5^{-3} = \frac{1}{5^3}\]

OpenStudy (anonymous):

so 1/2^3 x 2^1 ..now how do you solve?

jimthompson5910 (jim_thompson5910):

rewrite 2^1 as (2^1)/1 and then multiply straight across so 1*2^1 = 2 and 2^3 * 1 = 2^3 or 8 So you get 2/8 which reduces to 1/4

OpenStudy (anonymous):

so would 7^6 x 7^7 = 1/7?

jimthompson5910 (jim_thompson5910):

No, add the exponents

jimthompson5910 (jim_thompson5910):

what do you get